Hi,
My math teacher didn't do a good job explaining on how to slove a four consecutive integer problem can you help me?
Here's the problem:
Find four consecutive even integers such that four times the second is twenty more than the sum of the other three.

Answer by ewatrrr(24785)

 
Hi,
four consecutive even integers: x , x+2, x+4, x+6
4(x+2) = x + x+4 + x+ 6 + 20
4x + 8 = x + x+ 4 + x+ 6 + 20
x = 30-8
x = 22
four consecutive even integers: 22, 24, 26, 28
CHECKING our answer***
96 = 76 + 20
Wish You the Best in your Studies.